OpenStudy (anonymous):

solve and check the solution. -4(t-8)-(t+3)=3

OpenStudy (shane_b):

Do you know where to start?

OpenStudy (anonymous):

by combining like terms maybe?

OpenStudy (shane_b):

First, I would get rid of the parentheses...distribute the the -4 to to (t-8) and the negative sign to (t+3).

OpenStudy (shane_b):

Then the answer should become more obvious...add/subtract like terms and solve it from there.

OpenStudy (anonymous):

so it would be -4-32 and t-3?

OpenStudy (anonymous):

sorry a t in the -4

OpenStudy (shane_b):

-4t+32-t-3=3

OpenStudy (anonymous):

is the answer -7?

OpenStudy (shane_b):

-4t+32-t-3=3 -5t+29=3 -5t=-26 t=26/5

OpenStudy (anonymous):

ok I see where and how you got that answer. Thank you for your help.

OpenStudy (shane_b):

no problem...good luck :)